OpenStudy (anonymous):

Please help The radius of a circle is 15.2 cm. find the length of an arc of the circle intercepted by a central angle of 3pie/4

OpenStudy (anonymous):

I suppose they have given the angel in radian?

OpenStudy (anonymous):

well regardless of that, in order to find the arc length, the method will be the same whether the angle has be given in degrees or radians

OpenStudy (anonymous):

\[S = \theta*r\]

OpenStudy (anonymous):



OpenStudy (anonymous):

so in this case our s is unknown our θ = 3pi/4 and our radius = 15.2 cm

OpenStudy (anonymous):

\[S = \frac{ 3\Pi }{ 4 }* 15.2\] \[S = \frac{ 57 }{ 5 }\Pi \]

OpenStudy (anonymous):

which is basically 35.81 cm